OpenStudy (anonymous):

4x + 21/y - 17 For x = 9 and y = 26

OpenStudy (anonymous):

so 9 times 4 plus 21 divied by 26 -17

OpenStudy (anonymous):

9 times 4 is 36. add 21 to that which is 57 and 26 minus 17 is 9

OpenStudy (anonymous):

thsis iswhat got 36+21/26-17 36+21=57 26-17=9 36+21/26-17=57/9 57/9= 6.333333333333333

OpenStudy (anonymous):

now divide 57 by 9 and u get 6.33 repeating

OpenStudy (anonymous):